OpenStudy (anonymous):

Find the polynomial function with roots 11 and 2i

OpenStudy (nenadmatematika):

P(x)=(x-11)(x^2+4) Multiply and you'll get your solution......if you'd like some details....if x=2i is one root of the polynomial, then x=-2i is also the root, so polynomial can be divided by (x-2i)(x+2i)=x^2+4....for 11 I think it's clear :D